OpenStudy (anonymous):

What is the magnitude if the acceleration experienced by an electron in an E of 750 N/C ? How does the direction of the acceleration depende in the direction of the field at this point?

OpenStudy (anonymous):

F=m.a

OpenStudy (anonymous):

F=E.Q

OpenStudy (anonymous):

two formula?

OpenStudy (anonymous):

a=E.Q/m

OpenStudy (anonymous):

Q is the electron charge,and m is the mass of electron

OpenStudy (anonymous):

(750N/C) (9.1x10^-31) divided by ???

OpenStudy (anonymous):

direction?

OpenStudy (anonymous):

am i right with the Q ?

OpenStudy (anonymous):

no.. q=1.6*10^-19 m=9.1*10^-31

OpenStudy (anonymous):

then the e is 750 N/C ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what if -1.6*10^-19 ? what is the difference between this negative and the positive 1.6*10^-19 ?

OpenStudy (anonymous):

yeah...i forgot...both electron and proton have the same charge.. for electron...it's negative...and for proton it's positive.

OpenStudy (anonymous):

so for electron..-1.6*10^-19

OpenStudy (anonymous):

so i will use negative in this problem? because of electron?

OpenStudy (anonymous):

acceleration has opposite direction of field

OpenStudy (anonymous):

so you mean, even if it is electron, it has acceleration so it will be positive?

OpenStudy (anonymous):

the positive field will attract the electron

OpenStudy (anonymous):

oh okay that means, when it attracts the electron it will be positive ?

OpenStudy (anonymous):

how did u get the mass= 9.1*10^-31 ?

OpenStudy (anonymous):

use google

OpenStudy (anonymous):

is it constant? i cant find it on google. :( its shows Kinetic Energy and Electric Field

OpenStudy (anonymous):

wow thank you so much. God bless you :)